# Self-testing of genuine multipartite non-local and non-maximally entangled states

### v2

@@ Abstract
$\begin{split}P(x,y,z|\hat{x},\hat{y},\hat{z})&=\sum_{i,j,k}\lambda_{ijk}\operatorname{Tr}(\rho_{ijk}\Pi_{x|\hat{x}}\otimes\Pi_{y|\hat{y}}\otimes\Pi_{z|\hat{z}})\\ &\equiv\sum_{i,j,k}\lambda_{ijk}P_{ijk}(x,y,z|\hat{x},\hat{y},\hat{z})\end{split}$ (12)
where $\rho_{ijk}=\frac{\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k}\rho\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k}}{\lambda_{ijk}}$ is, at most, a three-qubit state. The coefficients $\lambda_{ijk}$ are determined as $\lambda_{ijk}=\operatorname{Tr}(\rho\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k})$, with $\lambda_{ijk}\geq 0$ $\forall i,j,k$ and $\sum_{i,j,k}\lambda_{ijk}=1$.
If the joint probability $P(x,y,z|\hat{x},\hat{y},\hat{z})$ meets the conditions outlined in 3, then it implies that the joint probability $P_{ijk}(x,y,z|\hat{x},\hat{y},\hat{z})$ will also adhere to the constraint equations. Consequently, the maximum attainable value of $\mathcal{S}_{3}$ across all three-qubit states denotes the optimal value achievable within tri-partite quantum states. ∎
Based on the findings presented in the preceding paragraphs, we are now ready to demonstrate the self-test of $|\Psi_{max}^{\star 3}\rangle_{ABC}$.
@@
Now we will provide self-testing of genuine $N$-partite non-local correlations. We will consider $\forall i,d_{i}=2$. So our conditions become:
@@
We have already seen that no local theory will satisfy $\mathcal{S}_{N}>0$. Let us consider the following general $N$-qubit pure state shared between $N$ parties $\mathcal{A}_{1},\mathcal{A}_{2},\dots,$ and $\mathcal{A}_{N}$:
@@
where $0<\alpha_{i}<\pi$ and $0\leq\phi_{i}<2\pi$. Now the state that will satisfy the equation 13 has amplitude of the following form:
@@
We will find the maximum value of $\mathcal{S}_{N}$ using equation 16. The result in [43] works for any $N$-party with two input and two output scenarios. Hence, our calculations for triparty case easily generalize for $N$-party also. Hence, we have self-testing of the state which gives the maximum value for $\mathcal{S}_{N}$.
@@
and projective measurements defined by equation 8. We performed numerical maximization of $\mathcal{S}_{3}^{\epsilon}$ across all states of the form 18 and measurement parameters, subject to the constraints outlined in equation 17. This process yielded $\mathcal{S}_{3}^{\epsilon}$ values for various noise parameter settings, ranging from 00 to $\frac{1}{3}$. Fig. 1 illustrates the maximum $\mathcal{S}_{3}^{\epsilon}$ values under $\mathcal{Q}_{3}$, three-qubit quantum states, and local model. Notably, we observed convergence between the maximum $\mathcal{S}_{3}^{\epsilon}$ values under $\mathcal{Q}_{3}$ and three-qubit quantum states, with an accuracy of approximately $10^{-6}$ within the error margin $0\leq\epsilon\leq 0.08$.
We have also calculated different entanglement monotones for these states against $GHZ$ and $W$ states in three and four parties, respectively. These results are depicted in 2. These states require very few resources to $GHZ$ and $W$ states.
@@
Ranendu Adhikary acknowledges funding and support from ISI DCSW Project No. PU/506/PL-MISC/521. We would like to acknowledge stimulating discussions with Kishor Bharti, Ashutosh Rai, Tamal Guha, Subhendu B. Ghosh, and Snehasish Roy Chowdhury.

### v3

@@ Abstract
$\begin{split}P(x,y,z|\hat{x},\hat{y},\hat{z})&=\sum_{i,j,k}\lambda_{ijk}\operatorname{Tr}[\rho_{ijk}(\Pi_{x|\hat{x}}\otimes\Pi_{y|\hat{y}}\otimes\Pi_{z|\hat{z}})]\\ &\equiv\sum_{i,j,k}\lambda_{ijk}P_{ijk}(x,y,z|\hat{x},\hat{y},\hat{z})\end{split}$ (12)
where $\rho_{ijk}=\frac{\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k}\rho\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k}}{\lambda_{ijk}}$ is, at most, a three-qubit state. The coefficients $\lambda_{ijk}$ are determined as $\lambda_{ijk}=\operatorname{Tr}[\rho(\Pi_{i}\otimes\Pi_{j}\otimes\Pi_{k})]$, with $\lambda_{ijk}\geq 0$ $\forall i,j,k$ and $\sum_{i,j,k}\lambda_{ijk}=1$.
If the joint probability $P(x,y,z|\hat{x},\hat{y},\hat{z})$ meets the conditions outlined in (3), then it implies that the joint probability $P_{ijk}(x,y,z|\hat{x},\hat{y},\hat{z})$ will also adhere to the constraint equations. Consequently, the maximum attainable value of $\mathcal{S}_{3}$ across all three-qubit states denotes the optimal value achievable within tri-partite quantum states.
@@
[…]
@@
We consider $p_{N}\equiv P(+,+,\dots,+|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})$ and $q_{N}\equiv P(-,-,\dots,-|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N})$. We have already seen that no local theory will satisfy $\mathcal{S}_{N}(:=p_{N}-q_{N})>0$. Let us consider the following general $N$-qubit pure state shared between $N$ parties $\mathcal{A}_{1},\mathcal{A}_{2},\dots,$ and $\mathcal{A}_{N}$:
@@
where $0<\alpha_{i}<\pi$ and $0\leq\phi_{i}<2\pi$. Now the state that will satisfy the equation (13) has amplitude of the following form:
@@
We will find the maximum value of $\mathcal{S}_{N}$ using equation (16) like we have already found the maximum value for $\mathcal{S}_{3}$ for triparty case. The result in [41] works for any $N$-party with two input and two output scenarios. Hence, our calculations for triparty case easily generalize for $N$-party also. Hence, we have self-testing of the state which gives the maximum value for $\mathcal{S}_{N}$.
@@
We performed numerical maximization of $\mathcal{S}_{3}^{\epsilon}$ across all states of the form 18 and measurement parameters, subject to the constraints outlined in equation (17). This process yielded $\mathcal{S}_{3}^{\epsilon}$ values for various noise parameter settings, ranging from 00 to $\frac{1}{3}$. Fig. 1 illustrates the maximum $\mathcal{S}_{3}^{\epsilon}$ values under $\mathcal{Q}_{3}$, three-qubit quantum states, and local model. Notably, we observed convergence between the maximum $\mathcal{S}_{3}^{\epsilon}$ values under $\mathcal{Q}_{3}$ and three-qubit quantum states, with an accuracy of approximately $10^{-6}$ within the error margin $0\leq\epsilon\leq 0.08$.
We have also calculated different entanglement monotones for these states against $GHZ$ and $W$ states in three and four parties, respectively. These results are depicted in Fig. 2. These states require very few resources to $GHZ$ and $W$ states.
@@
Ranendu Adhikary acknowledges funding and support from Indian Statistical Institute, Kolkata. We would like to acknowledge stimulating discussions with Kishor Bharti, Ashutosh Rai, Tamal Guha, Subhendu B. Ghosh, and Snehasish Roy Chowdhury.